# The category of anyon sectors for non-abelian quantum double models

### v1

@@ 3 Braided $\rm C^{*}$-tensor structure of $\mathbf{Amp}$ and $\mathbf{DHR}$
In Section 3.2 we use the assumption of Haag duality to endow $\mathbf{Amp}$ and $\mathbf{DHR}$ with braided ${\rm C}^{*}$-tensor structure. Most arguments in this section are straightforward adaptations of well-known constructions in the DHR superselection theory, see for example [SzlachanyiV93, NillS97, HalvorsonMueger, Ogata2021]. At this stage we do not know if the categories $\mathbf{Amp}_{f}$ and $\mathbf{DHR}_{f}$ are closed under the tensor product which we define for $\mathbf{Amp}$ and $\mathbf{DHR}$, a fact which will only be established in Proposition 7.7 and Lemma 7.9 of Section 7.
@@ 3.1 $\rm C^{*}$-structure, direct sums, and subobjects
Let us first remark that the categories $\mathbf{Amp}$, $\mathbf{Amp}_{f}$, $\mathbf{DHR}$, and $\mathbf{DHR}_{f}$ are $\rm C^{*}$-categories (see [GhezLimaRoberts] or  [neshveyev2013compact, Definition 2.1.1]). In this subsection we show that all these categories have finite direct sums and subobjects.
@@ 3.1.1 Direct sums and subobjects of amplimorphisms
We now establish the existence of subobjects.
@@
Proof : Consider the amplimorphism $\chi_{1}:{\mathcal{B}}\rightarrow M_{n}({\mathcal{B}})$ given by $\chi_{1}(O)=p\chi(O)p$. By Lemma 3.2 we have $p\in M_{n}({\mathcal{R}}(\Lambda))$ where $n$ is the degree of $\chi$, so $\chi_{1}$ is localized on $\Lambda$. Moreover, $p\chi_{1}(O)=p\chi(O)p=\chi(O)p$ and $\chi(\mathbbm{1})p=p=p\chi_{1}(\mathbbm{1})$ which shows that $p\in(\chi|\chi_{1})$.
The amplimorphism $\chi_{p}$ is also transportable. Indeed, let $\Lambda^{\prime}$ be some other cone. By transportability of $\chi$ there is an amplimorphism $\chi^{\prime}$ of degree $n^{\prime}$ localized on $\Lambda^{\prime}$ and an equivalence $U\in(\chi|\chi^{\prime})$. Consider the projection $q=U^{*}pU\in(\chi^{\prime}|\chi^{\prime})\subset M_{n^{\prime}}({\mathcal{R}}(\Lambda^{\prime}))$ and corresponding amplimorphism $\chi^{\prime}_{q}(O)=q\chi^{\prime}(O)q$ localized on $\Lambda^{\prime}$. Then
$pU\chi^{\prime}_{q}(O)=pUU^{*}pU\chi^{\prime}(O)U^{*}pU=p\chi(\mathbbm{1})p\chi(O)pU=p\chi(O)pU=\chi_{p}(O)pU$
and $pUU^{*}p=p\chi(\mathbbm{1})p=\chi_{p}(\mathbbm{1})$ while
@@
so $pU$ is an equivalence of $\chi_{p}$ and $\chi^{\prime}_{q}$.
The same construction yields a localized transportable amplimorphism $\chi_{2}$ corresponding to the orthogonal projector $q=\chi(\mathbbm{1})-p\in(\chi|\chi)$. One easily checks that the claim of the proposition are satisfies with $v=p$ and $w=q$.
@@ 3.1.2 Direct sums and subobjects in $\mathbf{DHR}$
The subcategory $\mathbf{DHR}$ is not closed under the direct sum described above, neither does the construction of subobjects stay in the $\mathbf{DHR}$ subcategory. However, $\mathbf{DHR}$ does have finite direct sums and subobjects, see [Ogata2021]. The subcategory $\mathbf{DHR}_{f}$ is closed under these direct sums, and any subobject of a finite endomorphism must again be finite, so that $\mathbf{DHR}_{f}$ also has finite direct sums and subobjects.
@@ 3.2 Braided $\rm C^{*}$-tensor structure of $\mathbf{Amp}$ and $\mathbf{DHR}$
Using the assumption of Haag duality for cones, we equip $\mathbf{Amp}$ and $\mathbf{DHR}$ with a monoidal product and a braiding, making them into braided $\rm{C}^{*}$-tensor categories (see Definition 2.1.1 of [neshveyev2013compact]). At this point it is not clear that the tensor product of two finite amplimorphisms, as defined below, is again finite (and in fact one can construct examples of irreducible anyon sectors whose monoidal product decomposes into infinitely many irreducibles, see for example [Fredenhagen94]). For this reason we can’t yet equip $\mathbf{Amp}_{f}$ and $\mathbf{DHR}_{f}$ with the structure of braided $\rm{C}^{*}$-tensor categories. It will be shown in Proposition 7.7 and Lemma 7.9 that $\mathbf{Amp}_{f}$ and $\mathbf{DHR}_{f}$ are in fact closed under the tensor product, and are therefore full braided $\rm C^{*}$-tensor subcategories of $\mathbf{Amp}$ and of $\mathbf{DHR}$ respectively.
@@ 3.2.1 Monoidal structure
The monoidal unit is the identity amplimorphism which is irreducible because ′′${\mathcal{B}}^{\prime\prime}={\mathcal{B}}({\mathcal{H}}_{0})$ since $\pi_{0}$ is irreducible and $\pi_{0}(\mathcal{A})\subset\mathcal{B}$. Since the monoidal product is strict, it is trivially compatible with the ${\rm C}^{*}$-structure. The subcategory $\mathbf{DHR}$ is closed under this monoidal product and contains the identity, it is therefore a monoidal subcategory of $\mathbf{Amp}$.
The monoidal product of objects is well defined thanks to Lemma 3.1 and the monoidal product of intertwiners is well defined thanks to Lemma 3.2.
@@ 3.2.2 Braiding
It is well known that the category of localized endomorphisms for models in two spatial dimensions can be given a braiding [frs1, frohlich1990braid]. Here we extend this to localized amplimorphisms.
@@
where $P:M_{n}(\mathbb{C})\otimes M_{m}(\mathbb{C})\to M_{m}(\mathbb{C})\otimes M_{n}(\mathbb{C})$ flips the tensor factors. Using standard arguments, on can check that indeed $\epsilon(\chi,\psi)\in(\psi\times\chi|\chi\times\psi)$, that $\epsilon(\chi,\psi)$ is independent of the choices of $\chi_{R},\psi_{L},U,V$, and that $\epsilon$ is indeed a braiding for $\mathbf{Amp}$. See for example [SzlachanyiV93, Prop. 5.2] for amplimorphisms, or [Naaijkens2011, Lemma 4.8], [Ogata2021, Definition 4.10], or [bols2024double, Lemma 2.9] for proofs of the analogous fact for the braiding of endomorphisms. This braiding restricts to the $\rm C^{*}$-tensor subcategory $\mathbf{DHR}$, so $\mathbf{DHR}$ is a braided $\rm C^{*}$-tensor subcategory of $\mathbf{Amp}$.
@@ 4.1 Reduction to unital amplimorphisms
We say $\widetilde{\Lambda}$ is slightly larger than $\Lambda$, denoted $\Lambda\Subset\widetilde{\Lambda}$, if there exists another cone $\Lambda^{\prime}\subset\widetilde{\Lambda}$ disjoint from $\Lambda$. That is, we can fit a cone in $\Lambda^{c}\cap\widetilde{\Lambda}$. The following Lemma is proven in exactly the same way as [Ogata2021, Lemma 5.11], and noting [KadisonR97, Corollary 6.3.5]. We include it here for the convenience of the reader, as we will use this result repeatedly.
@@
Let $\Lambda\Subset\widetilde{\Lambda}$ and let $p\in M_{n}({\mathcal{R}}(\Lambda))$ be an orthogonal projector. Then $p$ is infinite as a projector in $M_{n}({\mathcal{R}}(\widetilde{\Lambda}))$, and is Murray-von Neumann equivalent to $\mathbbm{1}_{n}$.
Proof : By assumption, there is a cone $\Lambda^{\prime}\subset\widetilde{\Lambda}$ that is disjoint from $\Lambda$. Since ${\mathcal{R}}(\Lambda^{\prime})$ is an infinite factor (see Sect. 2.2), so is $M_{n}({\mathcal{R}}(\Lambda^{\prime}))$ and we can apply the halving lemma [KadisonR97, Lemma 6.3.3] to find an isometry $V\in M_{n}({\mathcal{R}}(\Lambda^{\prime}))$ such that $VV^{*}<\mathbbm{1}_{n}$. Note that $V$ and $V^{*}$ commute with $p$ since they have disjoint supports. The map $x\mapsto xp$ for $x\in M_{n}({\mathcal{R}}(\Lambda^{\prime}))$ is a *-isomorphism from $M_{n}({\mathcal{R}}(\Lambda^{\prime}))$ onto $M_{n}({\mathcal{R}}(\Lambda^{\prime})p$ by [Kadison1983, Prop. 5.5.5]. In particular, this implies that $VV^{*}p\neq p$, and hence is a proper subprojection of $p$. Put $\widetilde{V}=pV$, then
@@
This shows that $p$ as a projection in $M_{n}({\mathcal{R}}(\widetilde{\Lambda}))$ is Murray von Neumann equivalent to its proper subprojection $pVV^{*}$ and thus $p$ is infinite in $M_{n}({\mathcal{R}}(\widetilde{\Lambda}))$. Murray-von Neumann equivalence to $\mathbbm{1}_{n}$ now follows immediately from Corollary 6.3.5 of [KadisonR97]. $\square$
@@
Proof : By Lemma 3.1 we have that the projector $\chi(\mathbbm{1})$ belongs to $M_{n}({\mathcal{R}}(\Lambda))$. By Lemma 4.1, it follows that $\chi(\mathbbm{1})$ is infinite as an element of $M_{n}({\mathcal{R}}(\widetilde{\Lambda}))$ and is Murray-von Neumann equivalent to $\mathbbm{1}_{n}\in M_{n}({\mathcal{R}}(\widetilde{\Lambda}))$. Therefore there exists an isometry $V\in M_{n}({\mathcal{R}}(\widetilde{\Lambda}))$ such that $VV^{*}=\chi(\mathbbm{1})$ and $V^{*}V=\mathbbm{1}_{n}$.
@@
Any localized and transportable amplimorphism $\chi$ is equivalent to a unital transportable amplimorphism $\chi^{\prime}$ localised in the same cone.
Proof : Let $\chi$ be localised in $\Lambda$. We have by transportability of $\chi$ that there exists an amplimorphism $\psi$ localised in a cone $\Lambda^{\prime}\Subset\Lambda$ such that $\psi\sim\chi$. We have by Lemma 4.2 that there exists a unital amplimorphism $\chi^{\prime}$ localised in $\Lambda$ such that $\chi^{\prime}\sim\psi$, so we have $\chi^{\prime}\sim\chi$. Transportability of $\chi^{\prime}$ is immediate by the transportability of $\chi$. $\square$
@@ 4.2 Proof of equivalence
For any cone $\Lambda$ and any $n\in{\mathbb{N}}$, fix a row vector $\mathbb{V}(\Lambda,n):=(V_{1},\cdots,V_{n})$ whose components are isometries $V_{i}\in{\mathcal{R}}(\Lambda)$ satisfying $V_{i}^{*}V_{j}=\delta_{ij}\mathbbm{1}$ and $\sum_{i=1}^{n}V_{i}V_{i}^{*}=\mathbbm{1}_{n}$. (Since ${\mathcal{R}}(\Lambda)$ is an infinite factor, we can repeatedly apply the halving lemma [KadisonR97, Lemma 6.3.3] to obtain such isometries). For any $\chi\in\mathbf{Amp}$ fix an allowed cone $\Lambda_{\chi}$ such that $\chi$ is localized on $\Lambda_{\chi}$ and write $\mathbb{V}_{\chi}=\mathbb{V}(\Lambda_{\chi},n)$, where $n$ is the degree of $\chi$.
@@
Even though we do not need it to prove Theorem 2.7, we can now easily obtain the following proposition.
@@ 5 Amplimorphisms from ribbon operators
In this section we construct for each half-infinite ribbon $\rho$ a full subcategory $\mathbf{Amp}_{\rho}$ of $\mathbf{Amp}$ whose objects are constructed as limits of certain ‘ribbon operators’ taking unitary representations of ${\mathcal{D}}(G)$ as input. (See Appendix B for the definition and basic properties of ribbons and ribbon operators). We will show in Section 7.1 that the categories $\mathbf{Amp}_{\rho}$ are equivalent to $\mathbf{Rep}_{f}\,{\mathcal{D}}(G)$ as braided $\rm C^{*}$-tensor categories, and in Section 7.2 that they are equivalent to the whole of $\mathbf{Amp}_{f}$, thus establishing the equivalence of $\mathbf{Amp}_{f}$ and $\mathbf{Rep}_{f}\,{\mathcal{D}}(G)$ as braided $\rm C^{*}$-tensor categories.

### v2

@@ 3 Braided $\rm C^{*}$-tensor structure of $\mathbf{Amp}$ and $\mathbf{DHR}$
[…]
@@ 3.1 $\rm C^{*}$-structure, direct sums, and subobjects
Let us first remark that the categories $\mathbf{Amp}$, $\mathbf{Amp}_{f}$, $\mathbf{DHR}$, and $\mathbf{DHR}_{f}$ are $\rm C^{*}$-categories (see [GLR85] or  [NT13, Definition 2.1.1]). In this subsection we show that all these categories have finite direct sums and subobjects.
@@ 3.1.1 Direct sums and subobjects of amplimorphisms
We now establish the existence of subobjects. Since at the moment we allow non-unital amplimorphisms, the construction is somewhat more elementary than the corresponding result for DHR endormorphisms (cf. [Oga22, Lemma 5.8]).
@@
Proof : Consider the amplimorphism $\chi_{1}:{\mathcal{B}}\rightarrow M_{n}({\mathcal{B}})$ given by $\chi_{1}(O):=p\chi(O)p$. By Lemma 3.2 we have $p\in M_{n}({\mathcal{R}}(\Lambda))$ where $n$ is the degree of $\chi$, so $\chi_{1}$ is localized on $\Lambda$. Moreover, $p\chi_{1}(O)=p\chi(O)p=\chi(O)p$ and $\chi(\mathbbm{1})p=p=p\chi_{1}(\mathbbm{1})$ which shows that $p\in(\chi|\chi_{1})$.
The amplimorphism $\chi_{1}$ is also transportable. Indeed, let $\Lambda^{\prime}$ be some other cone. By transportability of $\chi$ there is an amplimorphism $\chi^{\prime}$ of degree $n^{\prime}$ localized on $\Lambda^{\prime}$ and an equivalence $U\in(\chi|\chi^{\prime})$. Consider the projection $q=U^{*}pU\in(\chi^{\prime}|\chi^{\prime})\subset M_{n^{\prime}}({\mathcal{R}}(\Lambda^{\prime}))$ and corresponding amplimorphism $\chi^{\prime}_{q}(O):=q\chi^{\prime}(O)q$ localized on $\Lambda^{\prime}$. Then
$pU\chi^{\prime}_{q}(O)=pUU^{*}pU\chi^{\prime}(O)U^{*}pU=p\chi(\mathbbm{1})p\chi(O)pU=p\chi(O)pU=\chi_{1}(O)pU$
and $pUU^{*}p=p\chi(\mathbbm{1})p=\chi_{1}(\mathbbm{1})$ while
@@
so $pU$ is an equivalence of $\chi_{1}$ and $\chi^{\prime}_{q}$.
The same construction yields a localized transportable amplimorphism $\chi_{2}$ corresponding to the orthogonal projector $q=\chi(\mathbbm{1})-p\in(\chi|\chi)$. One easily checks that the claim of the proposition is satisfied with $v=p$ and $w=q$.
@@ 3.1.2 Direct sums and subobjects in $\mathbf{DHR}$
The subcategory $\mathbf{DHR}$ is not closed under the direct sum described above, neither does the construction of subobjects stay in the $\mathbf{DHR}$ subcategory. However, $\mathbf{DHR}$ does have finite direct sums and subobjects, see [Oga22]. The subcategory $\mathbf{DHR}_{f}$ is closed under these direct sums, and any subobject of a finite endomorphism must again be finite, so that $\mathbf{DHR}_{f}$ also has finite direct sums and subobjects.
@@ 3.2 Braided $\rm C^{*}$-tensor structure of $\mathbf{Amp}$ and $\mathbf{DHR}$
Using the assumption of Haag duality for cones, we equip $\mathbf{Amp}$ and $\mathbf{DHR}$ with a monoidal product and a braiding, making them into braided $\rm{C}^{*}$-tensor categories (see Definition 2.1.1 of [NT13]). At this point it is not clear that the tensor product of two finite amplimorphisms, as defined below, is again finite (and in fact one can construct examples of irreducible anyon sectors whose monoidal product decomposes into infinitely many irreducibles, see for example [Fre94]). For this reason we can’t yet equip $\mathbf{Amp}_{f}$ and $\mathbf{DHR}_{f}$ with the structure of braided $\rm{C}^{*}$-tensor categories. It will be shown in Proposition 7.7 and Lemma 7.9 that $\mathbf{Amp}_{f}$ and $\mathbf{DHR}_{f}$ are in fact closed under the tensor product, and are therefore full braided $\rm C^{*}$-tensor subcategories of $\mathbf{Amp}$ and of $\mathbf{DHR}$ respectively.
@@ 3.2.1 Monoidal structure
[…]
The monoidal product of objects is well defined thanks to Lemma 3.1 and the monoidal product of intertwiners is well defined thanks to Lemma 3.2. The monoidal product on $\mathbf{DHR}$ coincides with that defined in [Oga22] (see also the remarks around equations (1.28)–(1.29) there).
@@ 3.2.2 Braiding
It is well known that the category of localized endomorphisms for models in two spatial dimensions can be given a braiding [FRS89, FG90, Frö88]. Here we extend this to localized amplimorphisms.
@@
where $P:M_{n}(\mathbb{C})\otimes M_{m}(\mathbb{C})\to M_{m}(\mathbb{C})\otimes M_{n}(\mathbb{C})$ flips the tensor factors. Using standard arguments, on can check that indeed $\epsilon(\chi,\psi)\in(\psi\times\chi|\chi\times\psi)$, that $\epsilon(\chi,\psi)$ is independent of the choices of $\chi_{R},\psi_{L},U,V$, and that $\epsilon$ is indeed a braiding for $\mathbf{Amp}$. See for example [SV93, Prop. 5.2] for amplimorphisms, or [Naa11, Lemma 4.8], [Oga22, Definition 4.10], or [BKM24, Lemma 2.9] for proofs of the analogous fact for the braiding of endomorphisms. [^6] Note that in the case of approximate Haag duality (as in [Oga22]), one has to do some additional limiting procedure to define the braiding. This is because under the weaker localization properties, we do not necessarily have that $\rho\times\sigma=\sigma\times\rho$ if $\rho$ and $\sigma$ are approximately localized in disjoint cones. This braiding restricts to the $\rm C^{*}$-tensor subcategory $\mathbf{DHR}$, so $\mathbf{DHR}$ is a braided $\rm C^{*}$-tensor subcategory of $\mathbf{Amp}$.
@@ 4.1 Reduction to unital amplimorphisms
We say $\widetilde{\Lambda}$ is slightly larger than $\Lambda$, denoted $\Lambda\Subset\widetilde{\Lambda}$, if there exists another cone $\Lambda^{\prime}\subset\widetilde{\Lambda}$ disjoint from $\Lambda$. That is, we can fit a cone in $\Lambda^{c}\cap\widetilde{\Lambda}$. The following Lemma is proven in exactly the same way as [Oga22, Lemma 5.11], and noting [KR97a, Corollary 6.3.5]. We include it here for the convenience of the reader, as we will use this result repeatedly.
@@
Let $\Lambda\Subset\widetilde{\Lambda}$ and let $p\in M_{n}({\mathcal{R}}(\Lambda))$ be an orthogonal projector. Then $p$ is infinite as a projector in $)M_{n}({\mathcal{R}}(\widetilde{\Lambda}$)), and is Murray-von Neumann equivalent to $\mathbbm{1}_{n}$.
Proof : By assumption, there is a cone $\Lambda^{\prime}\subset\widetilde{\Lambda}$ that is disjoint from $\Lambda$. Since ${\mathcal{R}}(\Lambda^{\prime})$ is an infinite factor (see Sect. 2.2), so is $M_{n}({\mathcal{R}}(\Lambda^{\prime}))$ and we can apply the halving lemma [KR97a, Lemma 6.3.3] to find an isometry $V\in M_{n}({\mathcal{R}}(\Lambda^{\prime}))$ such that $VV^{*}<\mathbbm{1}_{n}$. Note that $V$ and $V^{*}$ commute with $p$ since they have disjoint supports. The map $x\mapsto xp$ for $x\in M_{n}({\mathcal{R}}(\Lambda^{\prime}))$ is a *-isomorphism from $M_{n}({\mathcal{R}}(\Lambda^{\prime}))$ onto $M_{n}({\mathcal{R}}(\Lambda^{\prime})p$ by [KR97, Prop. 5.5.5]. In particular, this implies that $VV^{*}p\neq p$, and hence is a proper subprojection of $p$. Put $\widetilde{V}=pV$, then
@@
This shows that $p$ as a projection in $)M_{n}({\mathcal{R}}(\widetilde{\Lambda}$)) is Murray von Neumann equivalent to its proper subprojection $pVV^{*}$ and thus $p$ is infinite in $)M_{n}({\mathcal{R}}(\widetilde{\Lambda}$)). Murray-von Neumann equivalence to $\mathbbm{1}_{n}$ now follows immediately from Corollary 6.3.5 of [KR97a]. $\square$
@@
Proof : By Lemma 3.1 we have that the projector $\chi(\mathbbm{1})$ belongs to $M_{n}({\mathcal{R}}(\Lambda))$. By Lemma 4.1, it follows that $\chi(\mathbbm{1})$ is infinite as an element of $)M_{n}({\mathcal{R}}(\widetilde{\Lambda}$)) and is Murray-von Neumann equivalent to $)\mathbbm{1}_{n}\in M_{n}({\mathcal{R}}(\widetilde{\Lambda}$)). Therefore there exists an isometry $)V\in M_{n}({\mathcal{R}}(\widetilde{\Lambda}$)) such that $VV^{*}=\chi(\mathbbm{1})$ and $V^{*}V=\mathbbm{1}_{n}$.
@@
Any localized and transportable amplimorphism $\chi$ is equivalent to a unital transportable amplimorphism $\chi^{\prime}$ localized in the same cone.
Proof : Let $\chi$ be localized in $\Lambda$. We have by transportability of $\chi$ that there exists an amplimorphism $\psi$ localized in a cone $\Lambda^{\prime}\Subset\Lambda$ such that $\psi\sim\chi$. We have by Lemma 4.2 that there exists a unital amplimorphism $\chi^{\prime}$ localized in $\Lambda$ such that $\chi^{\prime}\sim\psi$, so we have $\chi^{\prime}\sim\chi$. Transportability of $\chi^{\prime}$ is immediate by the transportability of $\chi$. $\square$
@@ 4.2 Proof of equivalence
We now show that instead of amplimorphisms, we can equivalently talk about endomorphisms. For any cone $\Lambda$ and any $n\in{\mathbb{N}}$, fix a row vector $\mathbb{V}(\Lambda,n):=(V_{1},\cdots,V_{n})$ whose components are isometries $V_{i}\in{\mathcal{R}}(\Lambda)$ satisfying $V_{i}^{*}V_{j}=\delta_{ij}\mathbbm{1}$ and $\sum_{i=1}^{n}V_{i}V_{i}^{*}=\mathbbm{1}_{n}$. (Since ${\mathcal{R}}(\Lambda)$ is an infinite factor, we can repeatedly apply the halving lemma [KR97a, Lemma 6.3.3] to obtain such isometries). For any $\chi\in\mathbf{Amp}$ fix an allowed cone $\Lambda_{\chi}$ such that $\chi$ is localized on $\Lambda_{\chi}$ and write $\mathbb{V}_{\chi}=\mathbb{V}(\Lambda_{\chi},n)$, where $n$ is the degree of $\chi$.
@@
Even though we do not need it to prove Theorem 2.7, we can now easily obtain the following proposition which says that the localized and transportable amplimorphisms are equivalent to the endomorphisms studied in [Oga22].
@@ 5 Amplimorphisms from ribbon operators
In this section we construct for each half-infinite ribbon $\rho$ a full subcategory $\mathbf{Amp}_{\rho}$ of $\mathbf{Amp}$ whose objects are constructed as limits of certain ‘ribbon operators’ taking unitary representations of ${\mathcal{D}}(G)$ as input. (See Appendix B for the definition and basic properties of ribbons and ribbon operators). From the equivalence of the localized and transportable amplimorphisms to DHR endomorphisms, this amounts to explicitly constructing examples of representations that satisfy the superselection criterion. More importantly, we can also define the intertwiners as (weak operator) limits of elements in the quasi-local algebra. In the notation of [Oga22], this amounts to finding explicit examples of the maps $T$ defined there, as well as how they act on the intertwiners.
The very concrete description of $\mathbf{Amp}_{\rho}$ and its intertwiners will allow us to identify the braiding and fusion in this category. We will use this to show in Section 7.1 that the categories $\mathbf{Amp}_{\rho}$ are equivalent to $\mathbf{Rep}_{f}\,{\mathcal{D}}(G)$ as braided $\rm C^{*}$-tensor categories, and in Section 7.2 that they are equivalent to the whole of $\mathbf{Amp}_{f}$, thus establishing the equivalence of $\mathbf{Amp}_{f}$ and $\mathbf{Rep}_{f}\,{\mathcal{D}}(G)$ as braided $\rm C^{*}$-tensor categories.